what is the slope of 4y-3x=8
Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

what is the slope of 4y-3x=8
=============
Solve for y --> the slope is the coefficient of x.
===============
eg,
what is the slope of 5y-2x = 3
5y = 2x + 3
y = (2/5)x + 3/5
The slope is 2/5
======================
3/5 is the y-intercept, BTW.

Answer by solver91311(24713)



Slope of the line that consists of the set of ordered pairs that is the solution set of







So the slope is:


John

My calculator said it, I believe it, that settles it